### v1

@@ 5.3 Local Clustering Coefficient
The local clustering coefficient of a node in a graph is used to quantify how close its neighbours are to being a clique (complete graph) i.e., how well connected the network is around a particular node. Mathematically, the local clustering coefficient ($LCC$) of a node $v\in G$ can be written as $LCC(v)=\frac{T(v)}{\delta(v)(\delta(v)-1)}$,(from  [45]). Here $\delta(v)$ denotes the degree of the node $v$. This metric can be easily calculated using mobile agents. $T(v)$ for a node $v$ can be calculated in $O(\Delta\log\lambda)$ rounds [Lemma 3]. The agent which already knows $\delta(v)$ at $v$ can now evaluate $LCC(v)$ using the formula.
@@ 6 Conclusion and Future Work
In this paper, we introduced algorithms for counting triangles via mobile agents. The agents enumerated the triangles based on nodes and edges. Finally, the total number of triangles in the graph was calculated. We used the triangle enumeration algorithm and applied it to the Truss Decomposition Problem, and to calculate the Triangle Centrality and Local Clustering Coefficient for each node.
It would interesting to establish lower bounds for the above computations in terms of both time complexity and memory per robot. Also, a lower bound on the number of robots used to execute the above algorithm can be an interesting investigation. Furthermore, the efficiency of the problem in the faulty robot model can be another task to consider in future.

### v2

@@ 5.3 Local Clustering Coefficient
The local clustering coefficient of a node in a graph is used to quantify how close its neighbours are to being a clique (complete graph) i.e., how well connected the network is around a particular node. Mathematically, the local clustering coefficient ($LCC$) of a node $v\in G$ can be written as $LCC(v)=\frac{T(v)}{\delta(v)(\delta(v)-1)}$,(from  [68]). Here $\delta(v)$ denotes the degree of the node $v$. This metric can be easily calculated using mobile agents. $T(v)$ for a node $v$ can be calculated in $O(\Delta\log\lambda)$ rounds [Lemma 4.3]. The agent that already knows $\delta(v)$ at $v$ can now evaluate $LCC(v)$ using the formula.
@@ 6 Conclusion and Future Work
In this paper, we introduce and analyse novel algorithms for triangle counting within the mobile agent model over an arbitrary anonymous graph. The agents first enumerate triangles based on the nodes and edges, and subsequently calculate the total number of triangles in the graph. We subsequently apply this triangle counting methodology to develop and analyse algorithms for solving the Truss Decomposition Problem and for computing important metrics such as Triangle Centrality and Local Clustering Coefficient for each node.
Future research could focus on several key areas to further enhance our approach. Establishing lower bounds for both time complexity and memory usage per agent would provide valuable insights into the efficiency and limitations of our algorithms. Additionally, exploring the lower bounds on the number of agents required to execute the algorithms could lead to more optimized and practical implementations. Another important direction is to investigate the performance of our algorithms in the presence of faulty agents, which would help in designing more robust solutions for real-world applications. Another interesting direction for future work could be to investigate scenarios where global parameters or their bounds, like the maximum degree $\Delta$ or the diameter $D$ of the graph, are not known in advance.